Calculate the firms return on equity as a percentage rounded to one decimal place if:

the risk-free rate is 2%, the return to the market is 7%, and the company's beta is 0.7
Answer #1

Expected return of overall market = Risk free rate + risk premium

= 2% + 7% = 9%(Answer)

Required rate of return of stock with beta b = 0.7 is:

= risk free rate + b * risk premium

=2 + 0.7*7 =2+4.9 = 6.9%(Answer)